OpenStudy (evansanchez):

Which expression has a value between –4 and –3? 1 - 2√10 4 - √15 3√5 - 7 √20 - 8

OpenStudy (mathmale):

Your best bet is probably to evaluate all four answer choices. Please show all of your results. Example: sqrt(20)-8=sqrt(4)sqrt(5)-8 = 2sqrt(5)-8 = about 4.47-8 = about -3.53.